# publications.bib

### v3

@@ 4.2. Perverse $t$-exactness
We are almost ready to deduce the result we need for torsion vanishing. To do this, we will first need to discuss when the additional assumptions of weak normalized regularity and $\mu$-regularity are superfluous, possibly under certain assumptions on $\ell$.
@@ 4.3. Verification of additional assumptions
We first need the following lemma which will allow us to base change to splitting fields.
@@
Let $G$ be a quasi-split connected reductive group with splitting field $E$. If $\phi_{T}$ is generic then $R\Gamma(W_{E},\tilde{\alpha}\circ\phi_{T}|_{W_{E}})$ is trivial for all absolute coroots $\tilde{\alpha}\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})$.
@@
We recall that, given a $\Gamma$-orbit of positive absolute coroots $\alpha\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}\backslash\Gamma$, if $E_{\alpha}$ denotes the reflex field of $\alpha$ then the representation of $\phantom{}{}^{L}T$ defined by $\alpha$ is given by choosing a representative $\tilde{\alpha}\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}$ of $\alpha$, and inducing the representation of $\hat{T}\rtimes W_{E_{\alpha}}/W_{E}$ defined by it to $W_{\mathbb{Q}_{p}}/W_{E}$. This reduces the claim to Schapiro’s Lemma. ∎
Other than the groups listed in Theorem 4.10, there are two more groups of interest to us. We will define them now.
@@
$G(\mathrm{SL}_{n,L}):=\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GL}_{n}\times_{\nu}\mathbb{G}_{m},$
$G(\mathrm{Sp}_{4,L}):=\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GL}_{n}\times_{\nu}\mathbb{G}_{m}.$
@@
    $G(\mathrm{SL}_{n,L})$.
If $\phi_{T}$ is a generic toral parameter for $G$ then $\phi_{T}$ is weakly normalized regular and regular. Moreover, for $(1)-(3)$, $\phi_{T}$ will be $\mu$-regular for all $\mu$, while, for $(4)$, $\phi_{T}$ will be $\mu$-regular for $\mu$ which are of the form $\prod_{\tau:L\hookrightarrow\overline{\mathbb{Q}}_{p}}\mu^{\prime}$ for $\mu^{\prime}$ a cocharacter of $\mathrm{GL}_{n}$.
@@
We establish weak normalized regularity, and suppress giving the proof that $\phi_{T}$ is regular as it is strictly easier. We note that since the condition of being generic is invariant under flipping the role of positive and negative roots, genericity implying that $\chi\otimes\delta_{B}^{1/2}\not\simeq\chi\otimes\delta_{B}^{-1/2}$ for $w\in W_{G}$ non-trivial implies also that $\chi^{w_{0}}\otimes\delta_{B}^{1/2}\not\simeq(\chi^{w_{0}}\otimes\delta_{B}^{-1/2})^{w}$, so it suffices to check the former.
For cases (1), (2), and (3), we may assume for simplicity that $L=\mathbb{Q}_{p}$ with the proof in general essentially being the same. If $G=\mathrm{GL}_{n}$ then this is [Ham2, Lemma 3.10].
We now consider the case of $G=\mathrm{U}_{n}$ defined with respect to a quadratic extension $E/\mathbb{Q}_{p}$. Suppose there exists a non-trivial $w\in W_{G}$ such that we have an isomorphism:
$\chi\otimes\delta_{B}^{1/2}\simeq(\chi\otimes\delta_{B}^{-1/2})^{w}$
of characters on $T(\mathbb{Q}_{p})$. We recall that $G_{E}\simeq\mathrm{GL}_{n,E}$ where $E/\mathbb{Q}_{p}$ denotes the quadratic extension defining the unitary group. By the definition of the modulus character in terms of the transformation character of Haar measures, we observe that the precomposition of $\delta_{B}$ with the Norm map $T(E)\rightarrow T(\mathbb{Q}_{p})$ gives the modulus character on the Borel of $\mathrm{GL}_{n,E}$. Therefore, by precomposing the previous isomorphism with this norm map, we obtain an analogous relationship of characters on the torus $T(E)$, which is the maximal torus of $\mathrm{GL}_{n,E}$. Then Lemma 4.21 reduces us to the $\mathrm{GL}_{n}$ case.
@@
For case (4), let $d=[L:\mathbb{Q}_{p}]$. Observe that we have an isomorphism $G(\mathrm{SL}_{n,L})_{L}\simeq H_{L}$, where
$H=\left\{(g_{i})\in\prod_{L\hookrightarrow\overline{\mathbb{Q}}_{p}}\mathrm{GL}_{n}:\mathrm{det}(g_{i})=\mathrm{det}(g_{j})\leavevmode\nobreak\ \forall i,j\right\}.$
Applying Lemma 4.21 again and arguing as for unitary groups, it suffices to work with $H_{L}$. We assume $L=\mathbb{Q}_{p}$ for notation simplicity. The maximal torus $T^{\prime}$ in $H$ can be identified with
@@
via the map $(t_{1},\dots,t_{d},t)\mapsto(\mathrm{diag}(t_{i},tt_{i}^{-1}))$. Since $W_{H}=\prod W_{\mathrm{GL}_{2}}$, consider any element $w^{\prime}\in W_{G}$, which we assume for notational simplicity is of the form
$(w,\dots,w,\mathrm{id},\dots,\mathrm{id}),$
where $w$ is the non-trivial element of the Weyl group of $\mathrm{GL}_{2}$, and we have $w$ in the first $k$ entries, for some integer $0\leq k\leq d$. The general case follows similarly. Observe that the isomorphism
$\chi\otimes\delta_{B}^{1/2}\simeq(\chi\otimes\delta_{B}^{-1/2})^{w^{\prime}}$
becomes
$\prod_{i=1}^{k}\chi_{i1}(t_{i}^{2}t^{-1})\chi_{i2}(tt_{i}^{-2})\simeq\prod_{i=k+1}^{d}|t_{i}^{-2}t|.$
If we substitute $t=x,t_{i}=x$ for $i=1,\dots,k$, while for $k+1\leq i\leq d$ we set $t_{i}=1$ if $i-k$ is odd and $t_{i}=x$ if $i-k$ is even, we see that we get $\prod_{i=1}^{k}\chi_{i1}(x)\chi_{i2}^{-1}(x)$ is isomorphic to either the trivial representation $\mathbf{1}$ or ||$\cdot$|| ⋅ |, which is a contradiction to genericity.
@@
We now show $\mu$-regularity. Again, for cases (1), (2), and (3), observe that if $G$ is of the form $\mathrm{Res}_{L/\mathbb{Q}_{p}}G^{\prime}$ and $T^{\prime}$ denotes the maximal torus of $G^{\prime}$ then we have an isomorphism
$\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})\simeq\prod_{\phi\in\mathrm{Hom}_{\mathbb{Q}_{p}}(L,\overline{L})}\mathbb{X}_{*}(T^{\prime}_{\overline{L}})$
@@
where $\overline{L}$ is an algebraic closure of $L$. Using this, we can without loss of generality assume that $L=\mathbb{Q}_{p}$. In the case that $G=\mathrm{GL}_{n},\mathrm{U}_{n}$, or $\mathrm{GU}_{n}$, this follows as in the proof of [Ham2, Corollary 10.16]. We recall briefly how this goes.
@@
One can consider the geometric dominant cocharacter $\mu=(1,0,\ldots,0,0)$ of $\mathrm{GL}_{n}$. This defines the standard representation $V_{\mathrm{std}}$ of $\hat{G}\simeq\mathrm{GL}_{n}$. This cocharacter is in particular minuscule so the weights form a closed Weyl group orbit with representative $(1,0,\ldots,0,0)$. From here, it easily follows that the difference of the weights appearing in $V_{\mathrm{std}}$ define coroots of $G$. In particular, it follows that, if $\phi_{T}$ is generic then it is strongly $\mu$-regular for $\mu=(1,0,\ldots,0)$ in the sense of Definition 4.13, and this implies the filtration on $T_{\mu}(\mathrm{nEis}(\mathcal{S}_{\phi_{T}}))$ splits by [Ham2, Theorem 10.10] for this $\mu$. Now, the tilting modules $\mathcal{T}_{\omega_{i}}=\Lambda^{i}(V_{\mathrm{std}})$ attached to the other fundamental coweights $\omega_{i}=(1^{i},0^{n-i})$ of $G$ can be realized as direct summands of $V_{\mathrm{std}}^{\otimes i}$, and it follows that $\phi_{T}$ is $\mu$-regular for $\mu=\omega_{i}$ by [Ham2, Proposition 10.12]. Since any dominant cocharacter can be written as a linear combination of fundamental weights, the claim for any $\mu$ now follows from [Ham2, Corollary 10.13]. The case of $\mathrm{GU}_{n}$ and $\mathrm{U}_{n}$ follows in a very similar way, using Lemma 4.21.
@@
For case (4), observe that as before, we can base change to $L$, and since $H$ is a subgroup of $\prod{\mathrm{GL}_{n}}$, all cocharacters $\mu$ of $H$ define products of cocharacters for $\mathrm{GL}_{n}$. Now, consider a cocharacter of the form $\mu=\prod_{\tau}\mu_{\tau}$, where $\mu_{\tau^{\prime}}=\mu_{\tau}$ for all $\tau,\tau^{\prime}$ and $\mu_{\tau}$ is a cocharacter of $\mathrm{GL}_{n}$. Note that every dominant minuscule cocharacter of $H$ will be of this form. This is because a cocharacter $\mu=\prod_{\tau}\mu_{\tau}$ of $\prod\mathrm{GL}_{n}$ factors through $H$ exactly when the composition with the determinant is equal for all $\tau$, and we see that for $\mu$ to be minuscule, $\mu_{\tau}$ must be one of the fundamental coweights $\omega_{i}$, which, after composing with the determinant, give different characters for $i\neq j$. Now, we observe that the same argument as above holds to show that the difference of Weyl conjugates define a coroot of $H$ for the cocharacter $\mu_{1}=\prod(1,0,\dots,0)$, while for all other cocharacters of the form $\mu=\prod\mu^{\prime}$, where $\mu^{\prime}$ is a fundamental weight of $\mathrm{GL}_{n}$, they appear as weights in some tensor power of the highest weight representation corresponding to $\mu_{1}$. The claim for any $\mu=\prod\mu^{\prime}$, where $\mu^{\prime}$ is a dominant cocharacter of $\mathrm{GL}_{n}$, follows by the same argument as above, using [Ham2, Corollary 10.13]. ∎
@@
Suppose moreover that $\ell\neq 2$ and $\ell$ is banal with respect to $L$ (i.e. $(q^{4}-1,l)=1$, where $q$ is the size of the residue field of $L$). If $\phi_{T}$ is a generic toral parameter for $G$ then $\phi_{T}$ is weakly normalized regular and regular. Moreover, for $(1)$, $\phi_{T}$ will be $\mu$-regular for all $\mu$, and, for $(2)$, $\phi_{T}$ will be $\mu$-regular for $\mu$ which are of the form $\prod_{\tau:L\hookrightarrow\overline{\mathbb{Q}}_{p}}\mu^{\prime}$ for $\mu^{\prime}$ a cocharacter of $\mathrm{GSp}_{4}$.
@@
We will first establish weak normalized regularity, and again suppress giving the proof that $\phi_{T}$ is regular as it is strictly easier. We recall that genericity implying that $\chi\otimes\delta_{B}^{1/2}\not\simeq\chi\otimes\delta_{B}^{-1/2}$ for $w\in W_{G}$ non-trivial implies also that $\chi^{w_{0}}\otimes\delta_{B}^{1/2}\not\simeq(\chi^{w_{0}}\otimes\delta_{B}^{-1/2})^{w}$, since being generic is stable under switching the roles of positive and negative roots, so it suffices to check the former. Again, for (1), we assume that $L=\mathbb{Q}_{p}$ for this part with the proof in general being more or less the same. We will show this by contradiction. Suppose on the contrary that there exists some $w\in W_{G}$ such that we have an isomorphism
(16) $\chi\otimes\delta_{B}^{1/2}\simeq(\chi\otimes\delta_{B}^{-1/2})^{w}.$
@@
$(t_{1},t_{2},t)\mapsto\begin{pmatrix}t_{1}&0&0&0\\ 0&t_{2}&0&0\\ 0&0&tt_{2}^{-1}&0\\ 0&0&0&tt_{1}^{-1}\end{pmatrix}$
as in [Tad, Page 135]. This allows us to write the character $\chi:T(\mathbb{Q}_{p})\rightarrow\overline{\mathbb{F}}_{\ell}^{*}$ as $\chi_{1}(t_{1})\chi_{2}(t_{2})\nu(t)$, for characters $\mathbb{Q}_{p}^{*}\rightarrow\overline{\mathbb{F}}_{\ell}^{*}$. Similarly, we can express the modulus character as
$\delta_{B}(t_{1},t_{2},t)=|t_{1}|^{4}|t_{2}|^{2}|t|^{-3}$
where ||$\cdot$|| ⋅ | is the norm character. We now check that (16) cannot hold for all seven non-trivial elements of the Weyl group.
@@
then evaluating equation (16) for this relationship reduces to
$\chi_{1}(t_{1})\chi_{2}(t_{2})\nu(t)|t_{1}|^{2}|t_{2}||t|^{-3/2}\simeq\chi_{1}(t_{1})\chi_{2}(t_{2}^{-1}t)\nu(t)|t_{1}|^{-2}|t_{2}||t|^{-1}|t|^{3/2}$
@@
$\chi_{2}(t)^{-1}\chi_{2}(t_{2})^{2}\simeq|t_{1}|^{-4}|t|^{2}$
so if we evaluate at $(t_{1},t_{2},t)=(x^{3},x^{2},x^{4})$ then we obtain
$\mathbf{1}\simeq|x|^{-4}$
which contradicts the assumption that $(p^{4}-1,\ell)=1$.
@@
$\chi_{1}(t_{1})\chi_{2}(t_{2})\nu(t)|t_{1}|^{2}|t_{2}||t|^{-3/2}\simeq\chi_{1}(t_{2})^{-1}\chi_{1}(t)\chi_{2}(t_{1})\nu(t)|t_{2}|^{2}|t|^{-2}|t_{1}|^{-1}|t|^{3/2}$
rearranging and cancelling terms we obtain
$\chi_{1}\chi_{2}^{-1}(t_{1})\chi_{2}\chi_{1}(t_{2})\chi_{1}(t)^{-1}\simeq|t_{1}|^{-3}|t_{2}||t|$
@@
$\chi_{1}^{-1}(x)\simeq|x|$
which contradicts genericity (See [Tad, Page 167] for the enumeration of $1$-parameter subgroups attached to the coroots in the parametrization (17). Note that we could also have substituted $(t_{1},t_{2},t)=(x,x,x)$ to obtain
$\chi_{1}(x)\simeq|x|^{-1}.$
@@
$\chi_{1}(x)^{-1}\simeq|x|$
which contradicts genericity. Note that we could also have substituted $(t_{1},t_{2},t)=(1,x,x)$ to obtain
$\chi_{1}(x)^{-1}\simeq|x|^{-1}.$
@@
$\chi_{2}\chi_{1}^{-1}(t_{2})\chi_{1}\chi_{2}(t_{1})\chi_{2}(t)^{-1}\simeq|t_{1}|^{-1}|t_{2}|^{-3}|t|^{2}$
so if we evaluate at $(t_{1},t_{2},t)=(x,1,x)$ then this gives
$\chi_{1}(x)\simeq|x|$
which contradicts genericity. Note that we could also have substituted $(t_{1},t_{2},t)=(1,x,x)$ to obtain
$\chi_{1}^{-1}(x)\simeq|x|^{-1}.$
@@
This concludes our discussion of weakly normalized regularity for $\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GSp}_{4}$.
@@
then we have $H_{L}\simeq G(\mathrm{Sp}_{4,L})_{L}$. Thus, we may reduce to the case of $H$. Since $H\subset\prod\mathrm{GSp}_{4}$, we may also use the parametrization in [Tad, 135] to see that the maximal torus $T^{\prime}$ is given by a parametrization

### v5

@@ 4.3. Perverse $t$-exactness
We are almost ready to deduce the result we need for torsion vanishing. To do this, we will first need to discuss when the additional assumptions of regularity and $\mu$-regularity are superfluous, possibly under certain assumptions on $\ell$.
@@ 4.4. Verification of additional assumptions
In this section, we will study how the conditions of regularity (Definition 4.15) and strong $\mu$-regularity (Definition 4.16) described in §4.3 are related to the condition of being generic in the particular case of the groups appearing in Theorem 4.13. In particular, we will see in many situations these extra conditions are in fact implied by the generic condition after imposing some additional constraints on $\ell$. In §4.4.1, we will study this for groups of type $A_{n}$, and then in §4.4.2 we will deal with the case of groups of type $C_{n}$. We then finally conclude with our main local results in §4.4.3 (Corollaries 4.27 and 4.29).
Before proceeding, we describe the following general Lemma which will allow us to base change to splitting fields.
@@
Let $G$ be a quasi-split connected reductive group with splitting field $F$. Suppose that $\ell\nmid[F:\mathbb{Q}]$ (e.g if $\ell$ is very decent with respect to $G$). Then $\phi_{T}$ is generic if and only if $R\Gamma(W_{F},\tilde{\alpha}\circ\phi_{T}|_{W_{F}})$ is trivial for all absolute coroots $\tilde{\alpha}\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})$.
@@
We recall that, given a $\Gamma$-orbit of absolute coroots $\alpha\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}/\Gamma$, if $F_{\alpha}$ denotes the reflex field of $\alpha$ then the representation of $\phantom{}{}^{L}T$ defined by $\alpha$ is given by choosing a representative $\tilde{\alpha}\in\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})^{+}$ of $\alpha$, and inducing the representation of $\hat{T}\rtimes W_{F_{\alpha}}/W_{F}$ defined by it to $W_{\mathbb{Q}_{p}}/W_{F}$. By Schapiro’s Lemma, we deduce that genericity is equivalent to $R\Gamma(W_{F_{\alpha}},\tilde{\alpha}\circ\phi_{T}|_{W_{F_{\alpha}}})\simeq 0$. Now, for all $i\in\mathbb{N}_{\geq 0}$, we recall that we have the maps defined by restriction and corestriction
$H^{i}(W_{F_{\alpha}},\tilde{\alpha}\circ\phi_{T}|_{W_{F_{\alpha}}})\xrightarrow{\mathrm{res}}H^{i}(W_{F},\tilde{\alpha}\circ\phi_{T}|_{W_{F}})\xrightarrow{\mathrm{cores}}H^{i}(W_{F_{\alpha}},\tilde{\alpha}\circ\phi_{T}|_{W_{F_{\alpha}}}),$
and that the composite is given by multiplication by $F:F_{\alpha}]$ (See [58, Chapter VIII,Proposition 4]). Therefore, since $\ell\nmid[F:F_{\alpha}]$ by our assumption, it follows that the vanishing of $R\Gamma(W_{F_{\alpha}},\tilde{\alpha}\circ\phi_{T}|_{W_{F_{\alpha}}})$ is equivalent to the vanishing of $R\Gamma(W_{F},\tilde{\alpha}\circ\phi_{T}|_{W_{F}})$, since multiplication by $F:F_{\alpha}]$ is invertible on $\overline{\mathbb{F}}_{\ell}$-vector spaces. This shows the desired claim. ∎
We will also want to consider a few additional groups other than those listed in Theorem 4.13. There are a few more groups of interest to us. We will define them now.
@@
$G(\mathrm{SL}_{n,L}):=\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GL}_{n}\times_{\nu,\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathbb{G}_{m}}\mathbb{G}_{m},$
$G(\mathrm{Sp}_{4,L}):=\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GSp}_{4}\times_{\nu,\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathbb{G}_{m}}\mathbb{G}_{m}.$
Now we will study the additional assumptions for the groups of type $A$.
4.4.1. Groups of type A
The first main result is of the following form.
@@ 4.4.1. Groups of type A
    $G(\mathrm{SL}_{2,L})$.
Suppose that $\ell\nmid 2[L:\mathbb{Q}_{p}]$ in cases (1)-(2) and that $\ell\nmid[L:\mathbb{Q}_{p}]$ in cases (3)-(4).
If $\phi_{T}$ is a generic toral parameter for $G$ then $\phi_{T}$ is also regular. Moreover, for (1)(3)(1)-(3), $\phi_{T}$ will be $\mu$-regular for all $\mu$, while, for $(4)$, $\phi_{T}$ will be $\mu$-regular for $\mu$ which are of the form $\prod_{\tau:L\hookrightarrow\overline{\mathbb{Q}}_{p}}\mu^{\prime}$ for $\mu^{\prime}$ a cocharacter of $\mathrm{GL}_{2}$; in particular, it will be $\mu$-regular for all minuscule $\mu$.
@@
We first establish the implication that $\phi_{T}$ generic implies $\phi_{T}$ is regular in cases (1)-(3). We may assume for simplicity that $L=\mathbb{Q}_{p}$ with the proof in general essentially being the same in light of Lemma 4.24 and our assumption on $\ell$. If $G=\mathrm{GL}_{n}$ then we write $\chi(t_{1},\ldots,t_{n})$ as $\prod_{i=1}^{n}\chi_{i}(t_{i})$, where $t_{i}$ are the coordinates of $T(\mathbb{Q}_{p})\simeq(\mathbb{Q}_{p}^{*})^{n}$. It then easily follows from evaluating the relationship $\chi\simeq\chi^{w}$ for some non-trivial $w$ at any coordinate of $T(\mathbb{Q}_{p})$ not fixed by $w$ and setting the remaining coordinates equal to $1$ to derive the relation $\chi_{i}\chi_{j}^{-1}\simeq\mathbf{1}$ for some $i\neq j$, which contradicts genericity. We now consider the case of $G=\mathrm{U}_{n}$ defined with respect to a quadratic extension $E/\mathbb{Q}_{p}$. Suppose there exists a non-trivial $w\in W_{G}$ such that we have an isomorphism:
$\chi\simeq\chi^{w}$
of characters on $T(\mathbb{Q}_{p})$. We recall that $G_{E}\simeq\mathrm{GL}_{n,E}$ where $E/\mathbb{Q}_{p}$ denotes the quadratic extension defining the unitary group. We then precompose the isomorphism $\chi\simeq\chi^{w}$ with the norm map $T(E)\rightarrow T(\mathbb{Q}_{p})$ to obtain an analogous relationship of characters on the torus $T(E)$, which is the maximal torus of $\mathrm{GL}_{n,E}$. Then Lemma 4.24 and our assumpition on $\ell$ reduces us to the $\mathrm{GL}_{n}$ case discussed above.
@@
We now turn to case (4), let $d=[L:\mathbb{Q}_{p}]$. Observe that we have an isomorphism $G(\mathrm{SL}_{2,L})_{L}\simeq H_{L}$, where
$H=\left\{(g_{i})\in\prod_{i=1}^{d}\mathrm{GL}_{2}:\mathrm{det}(g_{i})=\mathrm{det}(g_{j})\penalty 10000\ \forall i,j\right\}.$
where $d$ is the size of the set of embeddings $L\hookrightarrow\overline{\mathbb{Q}}_{p}$. Applying Lemma 4.24 again and arguing as for unitary groups, it suffices to work with $H_{L}$. The maximal torus $T^{\prime}$ in $H$ can be identified with
@@
via the map $(t_{1},\dots,t_{d},t)\mapsto(\mathrm{diag}(t_{i},tt_{i}^{-1}))$. We have $W_{H}=\prod_{i=1}^{d}\mathbb{Z}/2\mathbb{Z}$, where the non-trivial element act on each of the factors by sending $t_{i}\mapsto t_{i}^{-1}t$. Write $\chi(t_{1},\ldots,t_{d})=\chi_{1}(t_{1})\otimes\cdots\otimes\chi_{d}(t_{d})\otimes\nu(t)$ for characters $\chi_{1},\dots,\chi_{d},\nu$. If we have a permutation $w\in W_{H}$ which is non-trivial on the $i$th factor then setting $t_{i}=x$ and $t=x$ and setting the remaining coordinates $t_{j}$ for $i\neq j$ equal to $1$ we obtain the relationship
$\chi_{i}(t)\simeq\mathbf{1},$
which contradicts genericity.
@@
We now show $\mu$-regularity. As before, for cases (1), (2), and (3), observe that if $G$ is of the form $\mathrm{Res}_{L/\mathbb{Q}_{p}}G^{\prime}$ and $T^{\prime}$ denotes the maximal torus of $G^{\prime}$ then we have an isomorphism
$\mathbb{X}_{*}(T_{\overline{\mathbb{Q}}_{p}})\simeq\prod_{\phi\in\mathrm{Hom}_{\mathbb{Q}_{p}}(L,\overline{L})}\mathbb{X}_{*}(T^{\prime}_{\overline{L}}),$
@@
where $\overline{L}$ is an algebraic closure of $L$. Using this, we can without loss of generality assume that $L=\mathbb{Q}_{p}$. In the case that $G=\mathrm{GL}_{n},\mathrm{U}_{n}$, or $\mathrm{GU}_{n}$, this follows as in the proof of [24, Corollary 9.16]. We recall briefly how this goes.
@@
One can consider the geometric dominant cocharacter $\mu=(1,0,\ldots,0,0)$ of $\mathrm{GL}_{n}$. This defines the standard representation $V_{\mathrm{std}}$ of $\hat{G}\simeq\mathrm{GL}_{n}$. This cocharacter is in particular minuscule so the weights form a closed Weyl group orbit with representative $(1,0,\ldots,0,0)$. From here, it easily follows that the difference of the weights appearing in $V_{\mathrm{std}}$ define coroots of $G$. In particular, it follows that, if $\phi_{T}$ is generic then it is strongly $\mu$-regular for $\mu=(1,0,\ldots,0)$ in the sense of Definition 4.16, and this implies the filtration on $T_{\mu}(\mathrm{nEis}(\mathcal{S}_{\phi_{T}}))$ splits by [24, Theorem 9.10] for this $\mu$. Now, the tilting modules $\mathcal{T}_{\omega_{i}}=\Lambda^{i}(V_{\mathrm{std}})$ attached to the other fundamental coweights $\omega_{i}=(1^{i},0^{n-i})$ of $G$ can be realized as direct summands of $V_{\mathrm{std}}^{\otimes i}$ (See [24, Proposition 9.1,Theorem 9.2, and Proposition 9.3] for a review of how this works), and it follows that $\phi_{T}$ is $\mu$-regular for $\mu=\omega_{i}$ by [24, Corollary 9.12]. Since any dominant cocharacter can be written as a linear combination of fundamental coweights, the claim for any $\mu$ now follows from [24, Corollary 9.13]. The case of $\mathrm{GU}_{n}$ and $\mathrm{U}_{n}$ follows in a very similar way, using Lemma 4.24 and our assumption on $\ell$.
@@
For case (4), observe that as before, we can base change to $L$, and since $H$ is a subgroup of $\prod{\mathrm{GL}_{2}}$, all cocharacters $\mu$ of $H$ define products of cocharacters for $\mathrm{GL}_{2}$. Now, consider a cocharacter of the form $\mu=\prod_{\tau}\mu_{\tau}$, where $\mu_{\tau^{\prime}}=\mu_{\tau}$ for all $\tau,\tau^{\prime}$ and $\mu_{\tau}$ is a cocharacter of $\mathrm{GL}_{2}$. Note that every dominant minuscule cocharacter of $H$ will be of this form. This is because a cocharacter $\mu=\prod_{\tau}\mu_{\tau}$ of $\prod\mathrm{GL}_{2}$ factors through $H$ exactly when the composition with the determinant is equal for all $\tau$. Moreover, we see that for $\mu$ to be minuscule, $\mu_{\tau}$ must be one of the fundamental coweights $\omega_{i}$, and if some $\mu_{\tau}\neq\mu_{\tau^{\prime}}$, then after composing with the determinant we will get different characters. Now, we observe that the same argument as above holds to show that the difference of Weyl conjugates define a coroot of $H$ for the cocharacter $\mu_{1}=\prod(1,0)$, while for all other cocharacters of the form $\mu=\prod\mu^{\prime}$, where $\mu^{\prime}$ is a fundamental weight of $\mathrm{GL}_{2}$, they appear as weights in some tensor power of the highest weight representation corresponding to $\mu_{1}$. The claim for any $\mu=\prod\mu^{\prime}$, where $\mu^{\prime}$ is a dominant cocharacter of $\mathrm{GL}_{2}$, follows by the same argument as above, using [24, Corollary 9.13]. ∎
We now turn our attention to the groups of type $C$.
4.4.2. Groups of type C
The main lemma is the following.
@@ 4.4.2. Groups of type C
If $\phi_{T}$ is a generic toral parameter for $G$ then $\phi_{T}$ is regular. Moreover, in case $(1)$, the parameter $\phi_{T}$ will be $\mu$-regular for all $\mu$, and, for $(2)$, $\phi_{T}$ will be $\mu$-regular for $\mu$ which are of the form $\prod_{\tau:L\hookrightarrow\overline{\mathbb{Q}}_{p}}\mu^{\prime}$ for $\mu^{\prime}$ a cocharacter of $\mathrm{GSp}_{4}$; in particular, it will be $\mu$-regular for all minuscule $\mu$.
@@
We will first establish regularity, and suppress giving the proof that $\phi_{T}$ is regular as it is strictly easier. Again, for (1), we assume that $L=\mathbb{Q}_{p}$ for this part with the proof in general being essentially the same in light of Lemma 4.24. We will show this by contradiction. Suppose on the contrary that there exists some $w\in W_{G}$ such that we have an isomorphism
(20) $\chi\simeq\chi^{w}.$
@@
$(t_{1},t_{2},t)\mapsto\begin{pmatrix}t_{1}&0&0&0\\ 0&t_{2}&0&0\\ 0&0&tt_{2}^{-1}&0\\ 0&0&0&tt_{1}^{-1}\end{pmatrix}$,
as in [63, Page 135]. This allows us to write the character $\chi:T(\mathbb{Q}_{p})\rightarrow\overline{\mathbb{F}}_{\ell}^{*}$ as $\chi_{1}(t_{1})\chi_{2}(t_{2})\nu(t)$, for characters $\mathbb{Q}_{p}^{*}\rightarrow\overline{\mathbb{F}}_{\ell}^{*}$. We now check that (20) cannot hold for all seven non-trivial elements of the Weyl group.
@@
then evaluating equation (20) becomes
$\chi_{1}(t_{1})\chi_{2}(t_{2})\nu(t)\simeq\chi_{1}(t_{1})\chi_{2}(t_{2}^{-1}t)\nu(t)$
@@
$\chi_{2}(t)^{-1}\chi_{2}(t_{2})^{2}\simeq\mathbf{1}$
and setting $t=1$ this gives
$\chi_{2}^{-1}(t)\simeq\mathbf{1}$
which contradicts genericity (See [63, Page 167] for the enumeration of $1$-parameter subgroups attached to the coroots in the parametrization (21).
@@
$\chi_{1}(t_{1})\chi_{2}(t_{2})\nu(t)\simeq\chi_{1}(t_{2})^{-1}\chi_{1}(t)\chi_{2}(t_{1})\nu(t)$
rearranging and canceling terms we obtain
$\chi_{1}\chi_{2}^{-1}(t_{1})\chi_{2}\chi_{1}(t_{2})\chi_{1}(t)^{-1}\simeq\mathbf{1}$
@@
$\chi_{1}^{-1}(x)\simeq\mathbf{1}$
which contradicts genericity (See [63, Page 167] for the enumeration of $1$-parameter subgroups attached to the coroots in the parametrization (21).
@@
$\chi_{1}(x)^{-1}\simeq\mathbf{1}$
which contradicts genericity.
@@
$\chi_{2}\chi_{1}^{-1}(t_{2})\chi_{1}\chi_{2}(t_{1})\chi_{2}(t)^{-1}\simeq\mathbf{1}$
so if we evaluate at say $(t_{1},t_{2},t)=(x,1,x)$ then this gives
$\chi_{1}(x)\simeq\mathbf{1}$
which contradicts genericity.
@@
This concludes our discussion of generic regularity for $\mathrm{Res}_{L/\mathbb{Q}_{p}}\mathrm{GSp}_{4}$.
@@
then we have $H_{L}\simeq G(\mathrm{Sp}_{4,L})_{L}$. Thus, we may reduce to the case of $H$. Since $H\subset\prod\mathrm{GSp}_{4}$, we may also use the parametrization in [63, 135] to see that the maximal torus $T^{\prime}$ is given by a parametrization